find the value of k so that the system of the equation
x+y+3z=0
4x+3y+kz=0
2x+y+2z=0 has no a non trival solution

Answer by Edwin McCravy(20086)

find the value of k so that the system of the equation 

 x +  y + 3z = 0
4x + 3y + kz = 0
2x +  y + 2z = 0 

has non-trival solution

This is called a homogeneous system.  It has only zeros
on the right side.  The trivial solution for all 
homogeneous systems is (x,y,z) = (0,0,0).  But many
homogeneous systems have other, non-trivial solutions.

Make the augmented matrix:

[1 1 3 | 0]
[4 3 k | 0]
[2 1 2 | 0]

To get a 0 where the 4 is,
multiply row 1 temporarily
by -4 and add it to 1 times
row 2, then restore row 1: 

-4[1 1 3 | 0]
 1[4 3 k | 0]
  [2 1 2 | 0]

[1  1     3 | 0]
[0 -1 -12+k | 0]
[2  1     2 | 0]

To get a 0 where the 2 on
the bottom left is, multiply 
row 1 temporarily by -2 and 
add it to 1 times row 3, then 
restore row 1:

-2[1  1     3 | 0]
  [0 -1 -12+k | 0]
 1[2  1     2 | 0]

[1  1     3 | 0]
[0 -1 -12+k | 0]
[0 -1    -4 | 0]

To get a 0 where the -1 on
the 3rd row is, multiply row
2 temporarily by -1 and add 
it to 1 times row 3, then 
restore row 2:


  [1  1     3 | 0]
-1[0 -1 -12+k | 0]
 1[0 -1    -4 | 0]

[1  1     3 | 0]
[0 -1 -12+k | 0]
[0  0   8-k | 0]

Multiply row 2 by -1

[1  1     3 | 0]
[0  1  12-k | 0]
[0  0   8-k | 0]

The system will have
a non-trivial solution
if the bottom row has
only zeros, hence 8-k
must = 0.  So

8 - k = 0
    k = 8

That's the answer.
------------------------
However, if you want to 
find the non-trivial 
solutions, then substitute
k = 8 in the matrix

[1  1     3 | 0]
[0  1  12-k | 0]
[0  0   8-k | 0]

which becomes

[1  1      3 | 0]
[0  1 12-(8) | 0]
[0  0  8-(8) | 0]

[1  1  3 | 0]
[0  1  4 | 0]
[0  0  0 | 0]

Rewriting as a system of
equations:

1x + 1y + 3z = 0
0x + 1y + 4z = 0
0x + 0y + 0z = 0

or to show triangular form:

 x +  y + 3z = 0
      y + 4z = 0
          0z = 0

The 3rd equation tells us
that z can equal any arbitrary
number, so substitute z = a
in the 2nd equation

      y + 4z = 0
      y + 4a = 0
           y = -4a

Substitute z = a and y = -4a
into the first equation:

   x +  y +  3z = 0
 x + (-4a) + 3a = 0 
    x - 4a + 3a = 0
          x - a = 0
              x = a

So all non-trivial solutions
can be found from:

(x, y, z) = (a, -4a, a)

by choosing arbitrary values
for a.  The trivial solution
is when a = 0

(x, y, z) = (0, 0, 0)

One non-trivial solutions can be
found when a = 1

(x, y, z) = (1, -4, 1)

When a = 2, we get the non-trivial
solution

(x, y, z) = (2, -8, 2)

When a = -6.37, we get the non-trivial
solution

(x, y, z) = (-6.37, 25.48, -6.37)

etc., etc., etc.

Edwin